John Coote 1795-1881  Mary Holland 1799-1886 m George Coote  ? Spooner  James Bird Garnham  Mary Ann Coote  Frederick Coote  ? Clark  John Coote  ?  ? Woodrow Elizabeth Coote  Arthur Bentall Lake 1824-1889 Susannah Eleanor Coote 1824-1919  Audley Coote 1838-1915  Alice Aaron m  m  m  m  m  m  m Arthur Coote  John Coote  Julia Coote m  10 children 13 grandchildren 22 great-grandchildren 16 great-great-grandchildren  John Coote  ? Oliver  Daughter m John Oliver  12 children ( The Lake family tree)  ? Hansard  Daughter  Audley John Coote m  Lulu Coote

1